OpenStudy (gabylovesyou):

For a science experiment Corrine is adding hydrochloric acid to distilled water. The relationship between the amount of hydrochloric acid, x, and the amount of distilled water, y, is graphed below.

OpenStudy (gabylovesyou):

Which inequality best represents this graph? 2y − 3x < 0 3y − 2x < 0 2y − 3x > 0 3y − 2x > 0 @ganeshie8

OpenStudy (gabylovesyou):

@jhonyy9

OpenStudy (gabylovesyou):

@phi

OpenStudy (phi):

dashed line means < or > (as opposed to ≤ or ≥ ) but all your choices use < so that does not help

OpenStudy (gabylovesyou):

I got D... am i correct ?

OpenStudy (phi):

you can look at the dashed line and figure out is slope 4/6 = 2/3 the line goes through (0,0) . it 's equation is y - 0 = (2/3 x - 0 or just y = (2/3) x multiply both sides by 3 3y = 2x or 2x - 3y = 0

OpenStudy (phi):

or 3y - 2x =0 we see the area above the line is shaded. we want y > 2/3 x or 3y -2x > 0

OpenStudy (gabylovesyou):

37 - 2x > 0

OpenStudy (gabylovesyou):

3y****

OpenStudy (phi):

yes, that is what I got